OpenStudy (anonymous):

I need this ASAP: a flare is fired from the bottom of a gorge is visible only when the flare is above the rim. if it is fired with an initial velocity of 96ft/sec, and the gorge is 128ft deep, during what interval can the flare be seen?\[h=-16t^{2}+v_{o}t+h_{o}\]

OpenStudy (anonymous):

@Xishem

OpenStudy (anonymous):

initial height is 0 im assuming you have initial velocity substitute those in and find t when h = 128

OpenStudy (anonymous):

substitute the values where h = 128 and v= 96 h0= 0 u get a quadratic equation solve it and find the time

OpenStudy (anonymous):

your answer would lie between the 2 t values you calculate

OpenStudy (anonymous):

128 = -16t^2 +96t +0 16t^2 -96t + 128 = 0 use the quadratic formula and find the values, u will get two answers, subtract the small value from the large value

OpenStudy (anonymous):

\[128=16t^{2}+96t+0\]\[0=16t^{2}+96t-128\]\[0=t^{2}+6t-8\]\[0=(t+4)(t-2)\]\[t=-4 ~~~or~~~ t=2\]

OpenStudy (anonymous):

@thushananth01 ??

jimthompson5910 (jim_thompson5910):

That is correct, the flare is seen from t = 2 seconds to t = 4 seconds

OpenStudy (anonymous):

but the 4 is negative...?

jimthompson5910 (jim_thompson5910):

it shouldn't be, it's positive

jimthompson5910 (jim_thompson5910):

\[\Large t = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large t = \frac{-(-96)\pm\sqrt{(-96)^2-4(16)(128)}}{2(16)}\] \[\Large t = \frac{96\pm\sqrt{9216-(8192)}}{32}\] \[\Large t = \frac{96\pm\sqrt{1024}}{32}\] \[\Large t = \frac{96+\sqrt{1024}}{32} \ \text{or} \ t = \frac{96-\sqrt{1024}}{32}\] \[\Large t = \frac{96+32}{32} \ \text{or} \ t = \frac{96-32}{32}\] \[\Large t = \frac{128}{32} \ \text{or} \ t = \frac{64}{32}\] \[\Large t = 4 \ \text{or} \ t = 2\]

jimthompson5910 (jim_thompson5910):

Sorry for some reason, I didn't see that negative...not sure how you got -4, but it should be positive 4

OpenStudy (anonymous):

YES use the quadratic formula, my friend..u can find time directly..:)

OpenStudy (anonymous):

k thanks!!

OpenStudy (anonymous):

no probs:)